OpenStudy (anonymous):

Which logarithmic equation is equivalent to the exponential equation below 5^c=125 log5 125=c logc 125=5 log5c=125 log125c=5

OpenStudy (anonymous):

A xD

OpenStudy (anonymous):

@MarrBear STOP HELPING ME!!!

OpenStudy (anonymous):

Aw :c

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \log_AB^C ~~~~~⇒~~~~~A^C=B}\)

OpenStudy (cwrw238):

by definition if a^b = c then loga c = b

OpenStudy (anonymous):

@MarrBear LOL jk hahahha

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \rm \log_AB^C ~~~~~⇒~~~~~A^C=B}\) is also a proof for why there is no negative log, unless the base is negative, and the log is odd.

OpenStudy (cwrw238):

the important thing to remember is that the logarithm is the exponent to which the base is taken

OpenStudy (anonymous):

@cwrw238 @SolomonZelman @MarrBear Thank you all for the quick help I am extremely bad at logarithms

OpenStudy (solomonzelman):

Clarke, we can fix that if you want to :) • • • • You welcome !

OpenStudy (anonymous):

Remember the formula i gave you xD. its very simple

OpenStudy (anonymous):

@SolomonZelman @MarrBear I understand it is as easy as just plugging the logs into the formula but some dont fit, Like this one "log784=a"

OpenStudy (anonymous):

If it doesnt have a Base, the base is ALWAYS 10.

OpenStudy (anonymous):

@MarrBear i'll remember that

OpenStudy (anonymous):

So it would be 784= 10^a c:

OpenStudy (anonymous):

@MarrBear so log400=a would be 10^a=400

OpenStudy (anonymous):

Exactly! (:

OpenStudy (anonymous):

@MarrBear its so simple... -_-